OpenStudy (anonymous):

Rationalize the Denominator: √5² ------ 4-√6²

OpenStudy (anonymous):

\[\sqrt{6^{2}}=6\] so \[4-\sqrt{6^{2}}=4-6=-2\]

OpenStudy (anonymous):

hmm well with the other half of the problem i got my answer to be -5/2 because, i took √5² = 5 then √6² = 6 5 ---------- then took both to make 4-6 = -2 so -5/2?

OpenStudy (anonymous):

yes....the ans. will be -2.5

OpenStudy (anonymous):

-2.5?

OpenStudy (anonymous):

oh duh...ok got it